# Searching For Music Mixing Graphs: A Pruning Approach

### v1

@@ 1 Introduction
Refer to the following link for audio samples, pruned graphs, and appendices that contain additional details: https://sh-lee97.github.io/grafx-prune.
@@ 2 Differentiable Processing on Graphs
  • Mix — We ouput the sum of incomming signals.
@@ 2.1 Differentiable Implementation
    Except for the gain computation, its implementation is the same as the compressor.
@@
It is common to compute the graph output signal by processing each node one by one [15, 19]. However, this severely bottlenecks the computation speed for large mixing graphs. Therefore, we instead batch-process multiple nodes in parallel. For the graph in Figure 2b, we can batch-process $1$ equalizer e, $3$ noisegates n, and $5$ gain/pannings g sequentially. Then, we aggregate the intermediate outputs to $2$ subgroup mixes m (also in parallel). This part is identical to graph neural networks’ “message passing,” so we adopt their implementations [31]. We repeat these parallel computations until we reach the output node o. By doing so, we obtain the output faster; in this example, the number of sequential processing is reduced from $15$ (one-by-one) to $8$ (optimal). We empirically found that up to 5.8×5.8$\times$5.8 × speedup can be achieved for the pruned graphs with a RTX3090 GPU.
@@ 3.2 Candidate Sampling
      When the trial fails, we perform one of the followings. If $|V_{\mathrm{cand}}|>1$, we halve the candidate set, i.e., $r_{t}\leftarrow r_{t}/2$. Otherwise, i.e., if $|V_{\mathrm{cand}}|=1$, we finish the search of this type by removing it from the pool as $T_{\mathrm{pool}}\leftarrow T_{\mathrm{pool}}\setminus\{t\}$.
@@ 3.3 Optimization
We first train the console with $6$\mathrm{k}$$ steps. Then, we repeat $N_{\mathrm{iter}}=12$ rounds of pruning, each with $0.5$\mathrm{k}$$-step fine-tuning. As a result, the total number of optimization steps is the same as the previous console training. During the first $4$\mathrm{k}$$ steps of the pruning phase, we linearly increase the sparsity coefficient $\alpha_{\mathrm{p}}$ from 00 to $10^{-2}$. While we halved the full console optimization steps, which could lead to increased loss, it is justified due to the tight resource constraints. With a RTX3090 GPU, each song took about $56$\mathrm{m}$$, $29$\mathrm{m}$$, and $36$\mathrm{m}$$ using the brute-force, dry/wet, and hybrid methods, respectively.
@@ 3.4 Results
Finally, we analyze the effect of the value of tolerance $\tau$. Even with a very low tolerance $\tau=0.001$, we can nearly halve the number of processors, i.e., $\rho=0.49$. If we set the value too high, e.g., $\tau=0.1$, the resulting graphs are highly sparse but degrade their matches ($L_{\mathrm{a}}=0.499$, i.e., $0.090$ increase). The default setting of $\tau=0.01$ seems “just right,” balancing the match performance and graph sparsity. We can verify this with the spectrogram errors (bottom $3$ rows for each subplots; see Figure 7 and supplementary page). There is no noticeable degradation from the full consoles to $\tau=0.001$ and $0.01$.
@@
Finally, we pruned every song in the full dataset ensemble. Table 3 reports the results. The overall trend follows the evaluation subset results but with higher average audio loss ($0.509$ compared to the previous $0.422$). Figure 6 shows statistics of the $3$ datasets, initial mixing console graphs, and their pruned versions. MedleyDB has the smallest number of source tracks, an average of $17.6$. The Internal has the largest ($28.8$), closely followed by the MixingSecrets ($27.9$). The Internal dataset also has more subgroups, resulting in even larger mixing consoles. This is one potential cause of the higher sparsity of its pruned graphs; more processors were initially used to match the mix, and many of them were redundant. On average, $72.1$ processors ($108.5$ nodes) were remained for each song. Since each full mixing console has an average of $247.6$ processors ($280.1$ nodes), we achieved a pruning ratio of $0.692$.
@@ 4 Discussion
We started with a general formulation of retrieving mixing graphs from dry sources and mix. Then, we posed restrictions to cast the search to the pruning of mixing consoles, making it computationally feasible and obtaining more interpretable graphs. Next, with additional assumptions, we derived the iterative method that gradually removes negligible processors in a stochastic and greedy manner. As a result, instead of finding the exact optimal, our method gives (or “samples”) one of the close-to-optimal graphs. With the differentiable processors and relaxation of the pruning with the dry/wet weights, we optimized this objective via gradient descent. We explored $3$ methods to choose pruning candidates, comparing them in terms of their computational cost and graph sparsity. The hybrid method gave a good compromise, so we used it to gather over one thoudsand graph-audio pairs.
@@
  • The choice of processors and their implementations directly affect the match quality. Our setup, including the equalizer with zero-phase FIR and the reverb based on STFT mask, was motivated by its simplicity and fast computation on a GPU. However, other alternatives exist, e.g., parametric equalizer [20] and artificial reverberation [36], that allow more efficient computation in CPU and have compact parameterizations. Also, the spectrogram errors showed clear temporal patterns (vertical stripes), indicating that the loudness dynamics were not precisely matched. We suspect it is due to the ballistics approximation error, as recently reported [37]. If so, we might need a more sophisticated implementation of the compressor and noisegate. Also, the current method does not support time-varying parameters (or “automation”), which can cause audible errors. For example, we could not match fade-out, i.e., a gradual decrease in track loudness. Finally, we can add other processor types, e.g., saturation/distortion [38] or modulation effects [39].
  • We note several considerations to improve the current pruning method in terms of sparsity, match quality, and interpretability. First, we can modify the mixing console to reflect real-world practices more. For example, we can add send and return loops with additional processor chains. Post-equalizers for compressors and processors with multiple inputs or outputs (e.g., auxiliary sidechain and crossover filter) are also commonly used. Second, to prevent the pruning from harming the perceptual quality, the tolerance condition and the objective function must be appropriately designed. We used a simple multi-resolution STFT loss [32, 34], which has been reported to miss some perceptual features [40, 41]. Hence, we might need an alternative objective as a remedy [42]. Third, as discussed before, using average loss to determine the pruning might be inappropriate. Lastly, to increase the sparsity, more advanced neural network pruning techniques [23, 24] and domain-specific post-processing, e.g., merging LTI processors to a single processor with the combined effect, can be applied.
@@ A.1 Composition of audio processors
Most audio processors are designed to modify some specific properties of their input signals, e.g., magnitude response, loudness dynamics, and stereo width. As such, combining multiple processors is a common practice to achieve the full effect. Following the main text, we will use the terminology “graph” to represent this composition, although some previous works considered simple structures that allow a more compact form, e.g., a sequence. Now, we outline the previous attempts that tried to estimate the processing graph or its parameters from reference audio. These works differ in task and domain, processors, graph structure, and estimation methods. For example, if the references are dry sources and a wet mixture, this task becomes reverse engineering [11, 12, 13, 17]. In terms of the prediction targets, some fixed the graph and estimated only the parameters [7, 15, 17, 19, 44]. Others tried to predict the graph [13] or both [11, 12, 26, 16]. Table 4 summarizes and highlights such differences.
@@ A.2 Differentiable signal processing
Now, consider a composition of multiple differentiable processors; the entire graph remains differentiable due to the chain rule. However, the following practical considerations remain. If we fix the processing graph prior to the optimization and the graph size is relatively small, we can implement the “differentiable graph” following the existing implementations [19, 15]. That is, we compute every processor one by one in a pre-defined topological order. However, we have the following additional requirements. First, the pruning changes the graph during the optimization. Therefore, our implementation must take a graph and its parameters along with the source signals as input arguments for every forward pass. Note that this feature is also necessary when training a neural network that predicts the parameters of any given graph [11]. Second, the size of our graphs is much larger than the ones from previous works [19, 15, 26, 11]. In this case, the one-by-one computation severely bottlenecks the computation speed. Therefore, we derived a flexible and efficient graph computation algorithm (i) that can take different graphs for each forward pass as input and (ii) performs batched processing of multiple processors within a graph, utilizing the parallelism of GPUs. Finally, we note that other than the differentiation with respect to the input signals $\mathbf{S}$ and parameters $\mathbf{P}$, one might be interested in differentiation with respect to the graph structure $G$. The proposed pruning method performs this to a limited extent; deletion of a node $v_{i}$ is a binary operation that modifies the graph structure. We relaxed this to a continuous dry/wet weight $w_{i}$ and optimized it with the audio loss $L_{\mathrm{a}}$ and regularization $L_{\mathrm{p}}$.
@@ Appendix B Dry/wet Pruning Algorithm
Processors $13$ guitar effects, including non-linear processors, modulation effects, ambience effects, and equalizer filters.
@@
Graph Tree structure: applied a fixed chain of the $7$ processors for each track, and then summed the chain outputs altogether.

### v2

@@ 1 Introduction
Refer to the following link for audio samples, pruned graphs, code, and appendices with supplementary details: https://sh-lee97.github.io/grafx-prune.
@@ 2 Differentiable Processing on Graphs
  • Mix — We output the sum of the incomming signals.
@@ 2.1 Differentiable Implementation
    Except for the gain computation, its implementation is the same as that of the compressor.
@@
It is common to compute the graph output signal by processing each node one by one [15, 19]. However, this severely bottlenecks the computation speed for large mixing graphs. Therefore, we instead batch-process multiple nodes in parallel. For the graph in Figure 2b, we can batch-process $1$ equalizer e, $3$ noisegates n, and $5$ gain/pannings g sequentially. Then, we aggregate the intermediate outputs to $2$ subgroup mixes m (also in parallel). This part is identical to graph neural networks’ “message passing,” so we adopt their implementations [31]. We repeat these parallel computations until we reach the output node o. By doing so, we obtain the output faster; in this example, the number of sequential processing is reduced from $15$ (one-by-one) to $8$ (optimal). We empirically found that up to 5.8×5.8$\times$5.8 × speedup can be achieved for the pruned graphs with a RTX3090 GPU. Refer to our companion paper for further details [32].
@@ 3.2 Candidate Sampling
      When the trial fails, we perform one of the following. If $|V_{\mathrm{cand}}|>1$, we halve the candidate set, i.e., $r_{t}\leftarrow r_{t}/2$. Otherwise, i.e., if $|V_{\mathrm{cand}}|=1$, we finish the search of this type by removing it from the pool as $T_{\mathrm{pool}}\leftarrow T_{\mathrm{pool}}\setminus\{t\}$.
@@ 3.3 Optimization
We first train the console with $6$\mathrm{k}$$ steps. Then, we repeat $N_{\mathrm{iter}}=12$ rounds of pruning, each with $0.5$\mathrm{k}$$-step fine-tuning. As a result, the total number of optimization steps is the same as the previous console training. During the first $4$\mathrm{k}$$ steps of the pruning phase, we linearly increase the sparsity coefficient $\alpha_{\mathrm{p}}$ from 00 to $10^{-4}$. While we halved the full console optimization steps, which could lead to increased loss, it is justified due to the tight resource constraints. With a RTX3090 GPU, each song took about $56$\mathrm{m}$$, $29$\mathrm{m}$$, and $36$\mathrm{m}$$ using the brute-force, dry/wet, and hybrid methods, respectively.
@@ 3.4 Results
Finally, we analyze the effect of the value of tolerance $\tau$. Even with a very low tolerance $\tau=0.001$, we can nearly halve the number of processors, i.e., $\rho=0.49$. If we set the value too high, e.g., $\tau=0.1$, the resulting graphs are highly sparse but degrade their matches ($L_{\mathrm{a}}=0.499$, i.e., $0.090$ increase). The default setting of $\tau=0.01$ seems “just right,” balancing the match performance and graph sparsity. We can verify this with the spectrogram errors (bottom $3$ rows for each subplot; see Figure 7 and supplementary page). There is no noticeable degradation from the full consoles to $\tau=0.001$ and $0.01$.
@@
Finally, we pruned every song in the full dataset ensemble. Table 3 reports the results. The overall trend follows the evaluation subset results but with a higher mean audio loss ($0.509$ compared to the previous $0.422$). Figure 6 shows statistics of the $3$ datasets, initial mixing console graphs, and their pruned versions. MedleyDB has the smallest number of source tracks, an average of $17.6$. The Internal dataset has the largest ($28.8$), closely followed by the MixingSecrets ($27.9$). The Internal also has more subgroups, resulting in even larger mixing consoles. This is one potential cause of the higher sparsity of its pruned graphs; more processors were initially used to match the mix, and many of them were redundant. On average, $72.1$ processors ($108.5$ nodes) were remained for each song. Since each full mixing console has an average of $247.6$ processors ($280.1$ nodes), we achieved a pruning ratio of $0.692$.
@@ 4 Discussion
We started with a general formulation of retrieving mixing graphs from dry sources and mix. Then, we posed restrictions to restrict the search to the pruning of mixing consoles, making it computationally feasible and obtaining more interpretable graphs. Next, with additional assumptions, we derived the iterative method that gradually removes negligible processors in a stochastic and greedy manner. As a result, instead of finding the exact optimal, our method gives (or “samples”) one of the close-to-optimal graphs. With the differentiable processors and relaxation of the pruning with the dry/wet weights, we optimized this objective via gradient descent. We explored $3$ methods to choose pruning candidates, comparing them in terms of their computational cost and graph sparsity. The hybrid method gave a good compromise, so we used it to gather over one thousand graph-audio pairs.
@@
  • The choice of processors and their implementations directly affect the match quality. Our setup, including the equalizer with zero-phase FIR and the reverb based on STFT mask, was motivated by its simplicity and fast computation on a GPU. However, there are other options, such as parametric equalizer [20] and artificial reverberation [37], that allow more efficient computation in the CPU and have compact parameterizations. Also, the spectrogram errors showed clear temporal patterns (vertical stripes), indicating that the loudness dynamics were not precisely matched. We suspect it is due to the ballistics approximation error, as recently reported [38]. If so, we might need a more sophisticated implementation of the compressor and noisegate. In addition, the current method did not consider time-varying parameters (or “automation”), which can cause audible errors. For example, we could not match fade-out, i.e., a gradual decrease in track loudness. Finally, we can add other processor types, e.g., saturation/distortion [39] or modulation effects [40].
  • We note several considerations to improve the current pruning method in terms of sparsity, match quality, and interpretability. First, we can modify the mixing console to reflect real-world practices more. For example, we can add send and return loops with additional processor chains. Post-equalizers for compressors and processors with multiple inputs or outputs (e.g., auxiliary sidechain and crossover filter) are also commonly used. Second, to prevent the pruning from harming the perceptual quality, the tolerance condition and the objective function must be appropriately designed. We used a simple multi-resolution STFT loss [33, 35], which has been reported to miss some perceptual features [41, 42]. Therefore, we might need an alternative objective as a remedy [43]. Third, as discussed before, using average loss to determine the pruning might be inappropriate. Lastly, to increase the sparsity, more advanced neural network pruning techniques [23, 24] and domain-specific post-processing, e.g., merging LTI processors to a single processor with the combined effect, can be applied.
@@ A.1 Composition of audio processors
Most audio processors are designed to modify some specific properties of their input signals, e.g., magnitude response, loudness dynamics, and stereo width. As such, combining multiple processors is a common practice to achieve the full effect. Following the main text, we will use the terminology “graph” to represent this composition, although some previous works considered simple structures that allow a more compact form, e.g., a sequence. Now, we outline the previous attempts that tried to estimate the processing graph or its parameters from reference audio. These works differ in task and domain, processors, graph structure, and estimation methods. For example, if the references are dry sources and a wet mixture, this task becomes reverse engineering [11, 12, 13, 17]. In terms of the prediction targets, some fixed the graph and estimated only the parameters [7, 15, 17, 19, 45]. Others tried to predict the graph [13] or both [11, 12, 26, 16]. Table 4 summarizes such differences.
@@ A.2 Differentiable signal processing
Now, consider a composition of multiple differentiable processors; the entire graph remains differentiable due to the chain rule. However, the following practical considerations remain. If we fix the processing graph prior to the optimization and the graph size is relatively small, we can implement the “differentiable graph” following the existing implementations [19, 15]. That is, we compute every processor one by one in a pre-defined topological order. However, we have the following additional requirements. First, the pruning changes the graph during the optimization. Therefore, our implementation must take a graph and its parameters along with the source signals as input arguments for every forward pass. Note that this feature is also necessary when training a neural network that predicts the parameters of any given graph [11]. Second, the size of our graphs is much larger than the ones from previous works [19, 15, 26, 11]. In this case, the one-by-one computation severely bottlenecks the computation speed. Therefore, we derived a flexible and efficient graph computation algorithm (i) that can take different graphs for each forward pass as input and (ii) performs batched processing of multiple processors within a graph, utilizing the parallelism of GPUs. The details of this method are described in our companion paper [32]. Finally, we note that other than the differentiation with respect to the input signals $\mathbf{S}$ and parameters $\mathbf{P}$, one might be interested in differentiation with respect to the graph structure $G$. The proposed pruning method performs this to a limited extent; deletion of a node $v_{i}$ is a binary operation that modifies the graph structure. We relaxed this to a continuous dry/wet weight $w_{i}$ and optimized it with the audio loss $L_{\mathrm{a}}$ and regularization $L_{\mathrm{p}}$.
@@ Appendix B Dry/wet Pruning Algorithm
Processors $13$ guitar effects, including non-linear processors, modulation effects, ambiance effects, and equalizer filters.
@@
Graph Tree structure: applied a fixed chain of the $7$ processors for each track, and then summed the chain outputs together.